If total carrying cost is $35,000, carrying cost per unit is $5, and beginning inventory is 14,000 units, compute ending inventory. Round to two decimal places.
Answer #1

Solution:

Total inventory carrying cost = $35000

Carrying cost per unit = $5

Beginning inventory = 14000 units

Total number of inventory = total carrying cost/ carrying cost per unit

= 35000/5 = 7000 units

So, 7000 units represent the average inventory over a period of time.

So, 7000* 2 = 14000 units

This value is equal to beginning inventory.

Ending inventory = 14000 units – 14000 units

                    Ending inventory = 0 units.